### v1

# Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving

Abstract
Large language models (LLMs) have demonstrated transformative potential across various domains, yet they face significant challenges in knowledge integration and complex problem reasoning, often leading to hallucinations and unreliable outputs. Retrieval-Augmented Generation (RAG) has emerged as a promising solution to enhance LLMs accuracy by incorporating external knowledge. However, traditional RAG systems struggle with processing complex relational information and multi-step reasoning, limiting their effectiveness in advanced problem-solving tasks. To address these limitations, we propose CogGRAG, a cognition inspired graph-based RAG framework, designed to improve LLMs performance in Knowledge Graph Question Answering (KGQA). Inspired by the human cognitive process of decomposing complex problems and performing self-verification, our framework introduces a three-stage methodology: decomposition, retrieval, and reasoning with self-verification. By integrating these components, CogGRAG enhances the accuracy of LLMs in complex problem solving. We conduct systematic experiments with three LLM backbones on four benchmark datasets, where CogGRAG outperforms the baselines.
Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving
Yao Cheng, Yibo Zhao, Jiapeng Zhu, Yao Liu, Xing Sun, Xiang Li
1 Introduction
As a foundational technology for artificial general intelligence (AGI), large language models (LLMs) have achieved remarkable success in practical applications, demonstrating transformative potential across a wide range of domains Touvron et al. (2023); AI@Meta (2024); Yang et al. (2024). Their ability to process, generate, and reason with natural language has enabled significant advancements in areas such as machine translation Zhu et al. (2023), text summarization Basyal and Sanghvi (2023), and question answering Pan et al. (2023). Despite their impressive performance, LLMs still face significant limitations in knowledge integration beyond their pre-trained data boundaries. These limitations often lead to the generation of plausible but factually incorrect responses, a phenomenon commonly referred to as hallucinations, which undermines the reliability of LLMs in critical applications.
To mitigate hallucinations, Retrieval-Augmented Generation (RAG) Niu et al. (2023); Gao et al. (2023); Edge et al. (2024) has emerged as a promising paradigm, significantly improving the accuracy and reliability of LLM-generated contents through the integration of external knowledge. However, while RAG successfully mitigates certain aspects of hallucination, it still exhibits inherent limitations in processing complex relational information. As shown in Figure 1 (a), the core limitation of standard RAG systems lies in their reliance on vector-based similarity matching, which processes knowledge segments as isolated units without capturing their contextual interdependencies or semantic relationships Jin et al. (2024b). Consequently, traditional RAG implementations are inadequate for supporting advanced reasoning capabilities in LLMs, particularly in scenarios requiring complex problem-solving, multi-step inference, or sophisticated knowledge integration.
Figure 1: Representative workflow of two Retrieval-Augmented Generation paradigms for enhancing LLMs.
Recently, graph-based RAG Edge et al. (2024); Ma et al. (2024); Jin et al. (2024b); Mavromatis and Karypis (2024) has been proposed to address the limitations of conventional RAG systems by incorporating deep structural information from external knowledge sources. These approaches typically utilize knowledge graphs (KGs) to model complex relation patterns within external knowledge bases, employing structured triple representations (<entity, relation, entity>) to integrate fragmented information across multiple document segments. While graph-based RAG has shown promising results in mitigating hallucination and improving factual accuracy, several challenges remain unresolved:
$\bullet$ Complex Problem Reasoning. Complex problems cannot be resolved through simple queries; they typically require multi-step reasoning to derive the final answer. While previous work has sought to enhance reasoning capabilities through methods such as chain-of-thought (CoT) reasoning or multi-hop information retrieval, these search-based and iterative approaches still face inherent disadvantages Jin et al. (2024b); Mavromatis and Karypis (2024). As shown in Figure 1 (b), each step in the iterative process relies on the result of the previous step, indicating that errors occurred at previous steps can propagate to subsequent steps.
$\bullet$ Hallucinations. Despite the integration of external knowledge sources, large language models (LLMs) remain prone to generating inaccurate or fabricated responses when confronted with retrieval errors or insufficient knowledge coverage. However, prior research has lacked mechanisms for self-verification, which undermines the reliability of graph-based RAG in real-world applications.
To address these challenges, we propose CogGRAG, a Cognition inspired Graph RAG framework, designed to enhance the complex problem-solving capabilities of LLMs in Knowledge Graph Question Answering (KGQA) tasks. The framework introduces a three-stage methodology: (1) Decomposition, inspired by human problem-solving strategies, breaks down complex problems top-down into smaller, simpler sub-problems and forms a mind map. (2) Retrieval, built on all the decomposed sub-problems, this stage extracts multi-level structured information from external knowledge sources using knowledge graphs. This stage operates at both local and global levels: local retrieval identifies relevant information for individual sub-problems, while global retrieval establishes connections between multiple sub-problems, ensuring detailed and comprehensive knowledge integration. (3) Reasoning with Self-Verification, mimicking human self-reflection, this stage evaluates the reasoning process and verifies the accuracy of intermediate and final results. CogGRAG emulates human cognitive processes in tackling complex problems by employing a top-down decomposition strategy, comprehensive knowledge retrieval, and bottom-up reasoning with validation. By integrating these components, CogGRAG enhances the accuracy and reliability of LLM-generated content on KGQA tasks.
2 Related Work
Figure 2: The overall process of CogGRAG. Given a target question $Q$, CogGRAG first prompts the LLM to decompose it into a hierarchy of sub-problems in a top-down manner, constructing a structured mind map. Subsequently, CogGRAG prompts the LLM to extract both local and global key information from these sub-problems. Finally, CogGRAG guides the LLM to perform bottom-up reasoning and verification based on the mind map and the retrieved knowledge, until the final answer is derived.
Reasoning with LLM Prompting. Recent advancements in prompt engineering have demonstrated that state-of-the-art prompting techniques can significantly enhance the reasoning capabilities of LLMs on complex problems Wei et al. (2022); Yao et al. (2024); Besta et al. (2024). Chain of Thought (CoT Wei et al. (2022) explores how generating a chain of thought—a series of intermediate reasoning steps—significantly improves the ability of large language models to perform complex reasoning. Tree of Thoughts (ToT) Yao et al. (2024) introduces a new framework for language model inference that generalizes the popular Chain of Thought approach to prompting language models, enabling exploration of coherent units of text (thoughts) as intermediate steps toward problem-solving. The Graph of Thoughts (GoT) Besta et al. (2024) models the information generated by a LLM as an arbitrary graph, enabling LLM reasoning to more closely resemble human thinking or brain mechanisms. However, these methods remain constrained by the limitations of the model’s pre-trained knowledge base and are unable to address hallucination issues stemming from the lack of access to external, up-to-date information.
Knowledge Graph Augmented LLM. KGs Vrandečić and Krötzsch (2014) offer distinct advantages in enhancing LLMs with structured external knowledge. Early graph-based RAG methods Edge et al. (2024); He et al. (2024); Hu et al. (2024); Wu et al. (2023); Jin et al. (2024a) demonstrated this potential by retrieving and integrating structured, relevant information from external sources, enabling LLMs to achieve superior performance on knowledge-intensive tasks. However, these methods exhibit notable limitations when applied to complex problems. Recent advancements have introduced methods like chain-of-thought prompting to enhance LLMs’ reasoning on complex problems Sun et al. (2023); Ma et al. (2024); Jin et al. (2024b). Think-on-Graph Sun et al. (2023) introduces a new approach that enables the LLM to iteratively execute beam search on the KGs, discover the most promising reasoning paths, and return the most likely reasoning results. ToG-2 Ma et al. (2024) achieves deep and faithful reasoning in LLMs through an iterative knowledge retrieval process that involves collaboration between contexts and the KGs. GRAPH-COT Jin et al. (2024b) propose a simple and effective framework to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. However, iterative reasoning processes are prone to error propagation, as errors cannot be corrected once introduced.
3 Preliminaries
3.1 Problem Definition
In this paper, we focus on the KGQA task. The input of the task is the question $Q$, and the output is the answer $A$. Given the input question $Q$, we aim to design a graph-based RAG framework with an LLM backbone $p_{\theta}$ to enhance the complex problem-solving capabilities and generate the answer $A$.
3.2 Mind Map
[…]
4 Methods
CogGRAG implements the graph-based RAG paradigm, inspired by human cognitive processes. Specifically, it simulates human cognitive strategies by decomposing complex questions into a structured mind map and incorporates self-verification mechanisms during the reasoning process. The framework of CogGRAG consists of three key steps: decomposition, retrieval, and reasoning with self-verification. The overall workflow of the framework is illustrated in Figure 2.
4.1 Decomposition
When confronted with complex question, humans typically decompose them into a series of simpler sub-questions and address them incrementally. Inspired by this cognitive strategy, in this paper, we decompose the complex question $Q$ in a top-down manner, which are logically interconnected and form a structured mind map. The resulting mind map aids in identifying the necessary information required to answer, even if such information is not explicitly mentioned in the original question. For instance, in the question “The football manager who recruited David Beckham managed Manchester United during what timeframe?”, the key entity “Alex Ferguson” (the manager who recruited David Beckham) might be overlooked by traditional matching-based RAG methods, failing to derive the correct answer.
[…]
$\begin{split}&\left\{\right(q^{t+1}_{j},s^{t+1}_{j}),j=1,...,N\}\\ &=\text{Decompose}(q^{t},p_{\theta},prompt_{1}),\end{split}$ (1)
where $N$ is the number of sub-questions, which is adaptively determined by LLMs based on the logic of each sub-question. $prompt_{1}$ describes the question decomposition task in natural language, as shown in Table 6. After decomposition, each sub-question and its state are added as a new node $(q^{t+1}_{j},t+1,s^{t+1}_{j})$ at level $t+1$. Decomposition is terminated when all the states of the child nodes are “End”, constructing a mind map $M$.
Figure 3: The key information extraction with local and global level.
4.2 Retrieval
At the retrieval step, CogGRAG leverages the mind map generated from question decomposition to retrieve relevant information from an external KG. Firstly, to ensure detailed and comprehensive knowledge integration, we extract key information from all questions in the mind map from two distinct granularity levels to support subsequent retrieval processes. As shown in Figure 3, entities and triples are extracted from each individual question at the local level, while inter-question relationships are captured at the global level and represented as subgraphs. This process can be formalized as follows:
$Keys=\text{Extract}(M,p_{\theta},prompt_{2}),$ (2)
where $Keys$ consists of the sets of entities, triples and subgraphs extracted from the mind map $M$, guided by the extraction workflow outlined in $prompt_{2}$, as shown in Table 7 and Table 8.
Next, based on the extracted key information $Keys$, CogGRAG aims to retrieve relevant structured knowledge (triples) from the KG. Specifically, we expand each entity in the entity set of $Keys$ to its neighbors in the KG, forming triples that link the entity to its neighboring entities via relationships. These retrieved tuples are added to a candidate tuple set $\tilde{\mathcal{T}}$. Subsequently, to reduce redundant and noisy information, we employ the triples and the subgraphs in $Keys$ to prune the triples in $\tilde{\mathcal{T}}$ through similarity matching. In practice, only tuples in $\tilde{\mathcal{T}}$ with similarity scores exceeding a predefined threshold $\varepsilon$ are retained, resulting in the final triples set $\mathcal{T}$. This process can be formalized as follows:
$\begin{split}&\mathcal{T}=\\ &\Big{\{}x\in\tilde{\mathcal{T}}\mid\exists y\in Triple\cup Subgraph,\\ &sim(x,y)>\varepsilon\Big{\}},\end{split}$ (3)
where $Triple\in Keys$ represents the set of triples extracted from individual questions, $Subgraph\in Keys$ denotes the set of subgraphs extracted from inter-question relationships, $sim(\cdot)$ is implemented using cosine similarity, with $\varepsilon$ being a hyper-parameter. This pruning ensures that the tuples retrieved from the KG remain both concise and relevant, enhancing the accuracy of downstream reasoning tasks.
4.3 Reasoning with Self-Verification
During the inference phase, CogGRAG utilizes the triples $\mathcal{T}$ retrieved in the previous step to reason and generate answers for the input question. Drawing upon the self-verification theory Frederick (2005), which posits that humans engage in self-verification to validate their reasoning outcomes during the cognitive process, we design a dual-process framework inspired by dual-process theories Vaisey (2009). To this end, we incorporate two LLMs into this step: one dedicated to the reasoning process and the other to the self-verification of the reasoning results. This approach ensures a more robust and reliable inference mechanism by mimicking the human cognitive process of reasoning followed by critical evaluation.
Specifically, given the triples set $\mathcal{T}$ and the mind map $M$, we prompt the LLM to address all sub-questions in the mind map in a bottom-up manner until reaching the root node, which corresponds to the target question $Q$. At each level $t$, the reasoning model $LLM_{res}$ infers and generates an answer $a^{t}$ for the question $q^{t}$, where $a^{t}=LLM_{res}(\mathcal{T},p_{\theta},q^{t},\hat{M},prompt_{3})$ and $\hat{M}=\left\{(q^{T},\hat{a}^{T}),(q^{T-1},\hat{a}^{T-1}),...,(q^{t+1},\hat{a}^{t+1})\right\}$ comprises all the question-answer pairs that have already been successfully reasoned through. Subsequently, the verification model $LLM_{ver}$ reflects on the current reasoning results $a^{t}$, aiming to identify potential errors or inconsistencies in the reasoning process based on the accumulated reasoning path. If the verification model $LLM_{ver}$ detects an error at the current step $t$, it will prompt the reasoning model $LLM_{res}$ to re-think and generate a new answer $\hat{a}^{t}$, else $\hat{a}^{t}=a^{t}$. Notably, during reasoning, we explicitly prompt the LLM to respond with "I don’t know" for questions it cannot answer, rather than generating incorrect responses. This helps mitigate the hallucination problem commonly observed in LLMs during reasoning tasks. Ultimately, CogGRAG infers the answer $A$ to the target question $Q$, where $A=\hat{a}^{0}$. In summary, reasoning with self-verification employs a bottom-up reasoning process based on the mind map within a dual-process framework. All prompt templates used in CogGRAG can be found in the Appendix D.
Table 1: Datasets statistics.
Domain Dataset Data Split
Train Dev Test
Wikipedia HotpotQA 90564 7405 7405
WebQSP 3098 0 1639
CWQ 27,734 3480 3475
Academic GRBENCH-Academic 850
E-commerce GRBENCH-Amazon 200
Literature GRBENCH-Goodreads 240
Healthcare GRBENCH-Disease 270
Table 2: Overall results of our CogGRAG on three KBQA datasets. The best score on each dataset is highlighted.
Type Method HotpotQA CWQ WebQSP
RL EM RL EM RL EM
Without external knowledge graph
LLM-only Direct 19.1% 17.3% 31.4% 28.8% 51.4% 47.9%
CoT 23.3% 20.8% 35.1% 32.7% 55.2% 51.6%
With external knowledge graph
LLM+KG Direct+KG 27.5% 23.7% 39.7% 35.1% 52.5% 49.3%
CoT+KG 28.7% 25.4% 42.2% 37.6% 52.8% 48.1%
Graph-based RAG ToG 29.3% 26.4% 54.6% 49.1% 61.7% 57.4%
MindMap 27.9% 25.6% 46.7% 43.7% 56.6% 53.1%
Ours CogGRAG 34.4% 30.7% 56.3% 53.4% 59.8% 56.1%
5 Experiments
5.1 Experimental Settings
Datasets and evaluation metrics. In order to test CogGRAG’s complex problem-solving capabilities on KGQA tasks, we evaluate CogGRAG on three widely used complex KGQA benchmarks: (1) HotpotQA Yang et al. (2018), (2) WebQSP Yih et al. (2016), (3) CWQ Talmor and Berant (2018). Following previous work Li et al. (2023), full Wikidata Vrandečić and Krötzsch (2014) is used as structured knowledge sources for all of these datasets. Considering that Wikidata is commonly used for training LLMs, there is a need for a domain-specific QA dataset that is not exposed during the pretraining process of LLMs in order to better evaluate the performance. Thus, we also test CogGRAG on a recently released domain-specific dataset GRBENCH Jin et al. (2024b). All methods need interact with domain-specific graphs containing rich knowledge to solve the problem in this dataset. The statistics and details of these datasets can734 be found in Table 1. For all datasets, we use two evaluation metrics: (1) Exact match(EM): measures whether the predicted answer or result matches the target answer exactly. (2) Rouge-L(RL): measures the longest common subsequence of words between the responses and the ground truth answers.
Baselines. In our main results, we compare CogGRAG with three types state-of-the-art methods: (1) LLM-only methods without external knowledge, including direct reasoning and CoT Wei et al. (2022) by LLM. (2) LLM+KG methods integrate relevant knowledge retrieved from the KG into the LLM to assist in reasoning, including direct reasoning and CoT by LLM. (3) Graph-based RAG methods allow KGs and LLMs to work in tandem, complementing each other’s capabilities at each step of graph reasoning, including Mindmap Wen et al. (2023), Think-on-graph Sun et al. (2023) and Graph-CoT Jin et al. (2024b).
Experimental Setup. We conduct experiments with three LLM backbones: LLaMA2-13B Touvron et al. (2023), LLaMA3-8B AI@Meta (2024) and Qwen2.5-7B Yang et al. (2024). For all LLMs, we load the checkpoints from huggingface [^1] https://huggingface.co and use the models directly without fine-tuning. We implemented CogGRAG and conducted the experiments with one A800 GPU [^2] Our code is available at: https://anonymous.4open.science/r/RAG-5883. Consistent with the Think-on-graph settings, we set the temperature parameter to 0.4 during exploration and 0 during reasoning. The threshold $\epsilon$ in the retrieval step is set to $0.7$.
Due to the space limitation, we move details on datasets (Sec A), and baselines (Sec B) to appendices.
5.2 Main Results
We perform experiments to verify the effectiveness of our framework CogGRAG, and report the results in Table 2. We use Rouge-L (RL) and Exact match (EM) as metric for all three datasets. The backbone model for all the methods is LLaMA2-13B. From the table, the following observations can be derived: (1) CogGRAG demonstrates the best results in most cases. (2) Compared to methods that incorporate external knowledge, the LLM-only approach demonstrates significantly inferior performance. This performance gap arises from the lack of necessary knowledge in LLMs for reasoning tasks, highlighting the critical role of external knowledge integration in enhancing the reasoning capabilities of LLMs. (3) Graph-based RAG methods demonstrate superior performance compared to LLM+KG approaches. This performance advantage is particularly evident in complex problems, where not only external knowledge integration but also involving “thinking procedure” is essential. Methods ToG and MindMap exhibit enhanced performance by implementing iterative retrieval and reasoning processes on the KGs, thereby generating the most probable inference outcomes through “thinking” on the KGs.
We attribute CogGRAG ’s outstanding effectiveness in most cases primarily to its ability to decompose complex problems and construct a structured mind map prior to retrieval. This approach enables the construction of a comprehensive reasoning pathway, facilitating more precise and targeted retrieval of relevant information. Furthermore, CogGRAG incorporates a self-verification mechanism during the reasoning phase, further enhancing the accuracy and reliability of the final results. Together, these designs collectively enhance LLMs’ ability to tackle complex problems.
Table 3: Overall results of our CogGRAG with different backbone models on three KBQA datasets. We highlight the best score on each dataset in bold.
Type Method HotpotQA CWQ WebQSP
RL EM RL EM RL EM
LLM-only Qwen2.5-7B 15.3% 15.0% 25.4% 24.1% 46.7% 45.3%
LLaMA3-8B 17.5% 14.9% 30.3% 27.5% 50.4% 45.1%
LLaMA2-13B 19.1% 17.3% 31.4% 28.8% 51.4% 47.9%
LLM+KG Qwen2.5-7B+KG 24.2% 15.6% 33.8% 32.1% 46.7% 45.3%
LLaMA3-8B+KG 25.9% 21.4% 40.6% 35.3% 53.6% 49.1%
LLaMA2-13B+KG 27.5% 23.7% 39.7% 35.1% 52.5% 49.3%
Graph-Based RAG CogGRAG w/ Qwen2.5-7B 28.4% 27.1% 50.5% 45.7% 53.2% 51.6%
CogGRAG w/ LLaMA3-8B 32.1% 27.2% 53.5% 48.4% 57.2% 55.3%
CogGRAG w/ LLaMA2-13B 34.4% 30.7% 56.3% 53.4% 59.8% 56.1%
Table 4: Overall results of our CogGRAG on GRBENCH dataset. We highlight the best score on each dataset in bold.
Method E-commerce Literature Academic Healthcare
RL EM RL EM RL EM RL EM
LLaMA2-13B 7.1% 6.8% 5.4% 5.1% 5.4% 4.7% 4.3% 3.1%
Graph-CoT 26.4% 24.0% 26.7% 23.3% 19.3% 14.8% 28.1% 25.2%
CogGRAG 30.2% 28.7% 32.4% 30.1% 23.6% 21.5% 27.4% 25.6%
5.3 Performance with different backbone models
We evaluate how different backbone models affect its performance on three datasets HotpotQA, CWQ and WebQSP, and report the results in Table 3. We conduct experiments with three LLM backbones LLaMA2-13B, LLaMA3-8B and Qwen2.5-7B. For all LLMs, we load the checkpoints from huggingface and use the models directly without fine-tuning. From the table, we can observe the following key findings: (1) CogGRAG achieves the best results across all backbone models compared to the baseline approaches, demonstrating the robustness and stability of our method. (2) The performance of our method improves consistently as the model scale increases, reflecting enhanced reasoning capabilities. This trend suggests that our approach has significant potential for further exploration with larger-scale models, indicating promising scalability and adaptability.
5.4 Performance on domain-specific KG
Given the risk of data leakage due to Wikidata being used as pretraining corpora for LLMs, we further evaluate the performance of CogGRAG on a recently released domain-specific dataset GRBENCH Jin et al. (2024b) and report the results in Table 4. This dataset requires all questions to interact with a domain-specific KG. We use Rouge-L (RL) and Exact match (EM) as metrics for this dataset. The backbone model for all the methods is LLaMA2-13B Touvron et al. (2023). The table reveals the following observations: (1) CogGRAG continues to outperform in most cases. This demonstrates that our method consistently achieves stable and reliable results even on domain-specific knowledge graphs. (2) Both CogGRAG and Graph-CoT outperform LLaMA2-13B by more than 20%, which can be ascribed to the fact that LLMs are typically not trained on domain-specific data. In contrast, graph-based RAG methods can effectively supplement LLMs with external knowledge, thereby enhancing their reasoning capabilities. This result underscores the effectiveness of the RAG approach in bridging knowledge gaps and improving performance in specialized domains.
Figure 4: Ablation study on the main components of CogGRAG.
Table 5: Overall results of our CogGRAG on GRBENCH dataset. We highlight the best score on each dataset in bold.
Method Correct Missing Hallucination
LLaMA2-13B 19.1% 25.7% 55.2%
ToG 29.3% 20.2% 50.5%
MindMap 27.9% 22.4% 49.7%
CogGRAG 34.4% 40.6% 44.9%
5.5 Ablation Study
The ablation study is conducted to understand the importance of main components of CogGRAG. We select HotpotQA, CWQ and WebQSP as three representative datasets. First, we remove the problem decomposition module, directly extracting information for the target question, and referred to this variant as CogGRAG-nd (no decomposition). Next, we eliminate the global-level retrieval phase, naming this variant CogGRAG-ng (no global level). Finally, we remove the self-verification mechanism during the reasoning stage, designating this variant as CogGRAG-nv (no verification). These experiments aim to systematically assess the impact of each component on the overall performance of the framework. We compare CogGRAG with these three variants, and the results are presented in Figure 4. Our findings show that CogGRAG outperforms all the variants on the three datasets. Furthermore, the performance gap between CogGRAG and CogGRAG-nd highlights the importance of decomposition for complex problem reasoning in KGQA tasks.
5.6 Hallucination and Missing Evaluation
In the reasoning and self-verification phase of CogGRAG, we prompt the LLM to respond with “I don’t know” when encountering questions with insufficient or incomplete relevant information during the reasoning process. This approach is designed to mitigate the hallucination issue commonly observed in LLMs. To evaluate the effectiveness of this strategy, we test the model on the HotpotQA dataset, with results reported in Table 5. We categorize the responses into three types: “Correct” for accurate answers, “Missing” for cases where the model responds with “I don’t know,” and “Hallucination” for incorrect answers. As shown in the table results, our model demonstrates the ability to refrain from answering questions with insufficient information, significantly reducing the occurrence of hallucinations. This capability highlights the effectiveness of our approach in enhancing the reliability and truthfulness of the model’s responses.
6 Conclusion
In this paper, we proposed a graph-based RAG framework, CogGRAG, to enhance the complex reasoning capabilities of LLMs in KGQA tasks. Inspired by human cognitive processes when tackling complex problems, CogGRAG simulates human reasoning by decomposing complex questions into sub-questions and constructing a structured mind map. Furthermore, recognizing that humans often engage in self-verification during reasoning, CogGRAG incorporates a self-verification mechanism during the reasoning phase, leveraging a dual-process to perform reasoning followed by verification, thereby reducing the occurrence of hallucinations. By integrating decomposition, retrieval, and reasoning with self-verification, CogGRAG significantly improves the accuracy of LLMs in handling complex problems. Extensive experiments demonstrate that CogGRAG outperforms state-of-the-art baselines in the majority of cases.
Limitations
Although CogGRAG significantly enhances the reasoning capabilities of LLMs in complex problems on KGQA tasks, several limitations remain. First, while our method avoids iterative retrieval and reasoning, the processes of problem decomposition and self-verification inevitably introduce additional computational overhead. Thus, there is room for future improvements in optimizing reasoning efficiency for practical applications. Second, CogGRAG is currently limited to graph-based external knowledge and cannot be directly applied to document-based knowledge sources, which restricts its generalizability in real-world scenarios. In the future, we plan to explore methods that integrate multiple types of data structures as external knowledge to further enhance the reasoning capabilities of LLMs. Additionally, research on improving retrieval and reasoning efficiency remains a key focus for our ongoing work.
Ethical Considerations
Our research is fundamental in nature and is not directly linked to specific applications. Consequently, the potential for misuse or negative societal impacts depends on how others may apply our methodology. Furthermore, our work does not involve any stakeholders who may benefit or be disadvantaged, nor does it involve vulnerable groups. All datasets used in this study are publicly available and widely utilized in the research community, ensuring no privacy risks and alignment with their intended purpose for scientific inquiry.
Appendix A Datastes
Here, we introduce the four datasets used in our experiments in detail. For HotpotQA and CWQ datasets, we evaluated the performance of all methods on the dev set. For WebQSP dataset, we evaluated the performance of all methods on the train set. The statistics and details of these datasets can be found in Table 1, and we describe each dataset in detail below:
HotpotQA is a large-scale question answering dataset aimed at facilitating the development of QA systems capable of performing explainable, multi-hop reasoning over diverse natural language. It contains 113k question-answer pairs that were collected through crowdsourcing based on Wikipedia articles.
WebQSP is a dataset designed specifically for question answering and information retrieval tasks, aiming to promote research on multi-hop reasoning and web-based question answering techniques. It contains 4,737 natural language questions that are answerable using a subset Freebase KG Bollacker et al. (2008).
CWQ is a dataset specially designed to evaluate the performance of models in complex question answering tasks. It is generated from WebQSP by extending the question entities or adding constraints to answers, in order to construct more complex multi-hop questions.
GRBENCH is a dataset to evaluate how effectively LLMs can interact with domain-specific graphs containing rich knowledge to solve the desired problem. GRBENCH contains 10 graphs from 5 general domains (academia, e-commerce, literature, healthcare, and legal). Each data sample in GRBENCH is a question-answer pair.
Appendix B Baselines
In this subsection, we introduce the baselines used in our experiments, including LLaMA2-13B, LLaMA3-8B, Qwen2.5-7B, Chain-of-Thought (CoT) prompting, Think-on-graph (ToG), MindMap and Graph-CoT.
LLaMA2-13B Touvron et al. (2023) is a member of the LLM series developed by Meta and is the second generation version of the LLaMA (Large Language Model Meta AI) model. LLaMA2-13B contains 13 billion parameters and is a medium-sized large language model.
LLaMA3-8B AI@Meta (2024) is an efficient and lightweight LLM with 8 billion parameters, designed to provide high-performance natural language processing capabilities while reducing computing resource requirements.
Qwen2.5-7B Yang et al. (2024) is an efficient and lightweight LLM launched by Alibaba with 7 billion parameters. It aims at making it more efficient, specialized, and optimized for particular applications.
Chain-of-Thought Wei et al. (2022) is a method designed to enhance a model’s reasoning ability, particularly for complex tasks that require multiple reasoning steps. It works by prompting the model to generate intermediate reasoning steps rather than just providing the final answer, thereby improving the model’s ability to reason through complex problems.
Think-on-graph Sun et al. (2023) proposed a new LLM-KG integration paradigm, “LLM $\otimes$ KG”, which treats the LLM as an agent to interactively explore relevant entities and relationships on the KG and perform reasoning based on the retrieved knowledge. It further implements this paradigm by introducing a novel approach called “Graph-based Thinking”, in which the LLM agent iteratively performs beam search on the KG, discovers the most promising reasoning paths, and returns the most likely reasoning result.
MindMap Wen et al. (2023) propose a novel prompting pipeline, that leverages KGs to enhance LLMs’ inference and transparency. It enables LLMs to comprehend KG inputs and infer with a combination of implicit and external knowledge. Moreover, it elicits the mind map of LLMs, which reveals their reasoning pathways based on the ontology of knowledge.
Graph-CoT Jin et al. (2024b) propose a simple and effective framework called Graph Chain-of-thought to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. Moreover, it manually construct a Graph Reasoning Benchmark dataset called GRBENCH, containing 1,740 questions that can be answered with the knowledge from 10 domain graphs.
Figure 5: Case of CogGRAG.
Appendix C Case Studies
In this section, we present a case analysis of the HotpotQA dataset to evaluate the performance of CogGRAG. As illustrated in Figure 5, CogGRAG decomposes the input question into two logically related sub-questions. Specifically, the complex question is broken down into “In which league cup” and “What was the sponsored name”, allowing the system to first identify the league cup and then determine its sponsored name based on the identified league. These sub-questions form a mind map, capturing the relationships between different levels of the problem. Next, CogGRAG extracts key information from all sub-questions, including both local level information within individual sub-questions and global level information across different sub-questions. A subgraph is constructed to represent the interconnected triples within the subgraph, enabling a global perspective to model the relationships between different questions. The retrieved triples are pruned based on similarity metrics. All information retrieved from the KG is represented in the form of triples. Using this knowledge, CogGRAG prompts the LLM to perform bottom-up reasoning and self-verification based on the constructed mind map. Through this process, the system ultimately derives the answer to the target question: “Carabao Cup.” This case demonstrates the effectiveness of CogGRAG in handling complex, multi-step reasoning tasks by leveraging hierarchical decomposition, structured knowledge retrieval, and self-verification mechanisms.
Additionally, Figure 6, Figure 8 and Figure 7 illustrate the process of prompting the large language model (LLM) to perform reasoning and self-verification, which provides a detailed breakdown of how the model generates and validates intermediate reasoning steps, ensuring the accuracy and reliability of the final output.
Figure 6: The prompt case of reasoning.
Figure 7: The prompt case of reasoning with the completed reasoning.
Figure 8: The prompt case of self-verification.
Appendix D Prompts in CogGRAG
In this section, we show all the prompts that need to be used in the main experiments as shown in Table 6, Table 7, Table 8, Table 9, Table 10 and Table 11.
Table 6: The prompt template for decomposition.
Decomposition prompt
Prompt head: “Your task is to decompose the given question Q into sub-questions.
You should based on the specific logic of the question to determine the number of sub-questions
and output them sequentially. ”
Instruction: “Please only output the decomposed sub-questions as a string in list format, where each
element represents the text of a sub-question, in the form of ’[\"subq1\", \"subq2\", \"subq3\"]’.
For each sub-questions, if you consider the sub-question to be sufficiently simple and no
further decomposition is needed, then output \"End.\", otherwise, output \"Continue.\".
Please strictly follow the format of the example below when answering the question.
Here are some examples: ”
“ Input: “What year did Guns N Roses perform a promo for a movie starring Arnold Schwarzenegger
as a former New York Police detective?”
Output: [ { “Sub-question”: “What movie starring Arnold Schwarzenegger as a former New York Police
detective is being referred to?”,
“State”: “Continue.” },
{ “Sub-question”: “In what year did Guns N Roses perform a promo for the movie mentioned
in sub-question #1?”,
“State”: “End.” } ]
Input: “What is the name of the fight song of the university whose main campus is in Lawrence, Kansas
and whose branch campuses are in the Kansas City metropolitan area?”
Output: [ { “Sub-question”: “Which university has its main campus in Lawrence, Kansas and branch
campuses in the Kansas City metropolitan area?”,
“State”: “End.”
{ “Sub-question”: “What is the name of the fight song of the university identified in sub-question #1?”,
“State”: “End.” } ]
Input: “Are the Laleli Mosque and Esma Sultan Mansion located in the same neighborhood?”
Output: [ { “Sub-question”: “Where is the Laleli Mosque located?”,
“State”: “End.”
{ “Sub-question”: “Where is the Esma Sultan Mansion located?”,
“State”: “End.” },
{ “Sub-question”: “Are the locations of the Laleli Mosque and the Esma Sultan Mansion in the same
neighborhood?”,
“State”: “End.” } ] ”
Input: Question Q
Output:
Table 7: The prompt template for extraction on local level.
Extraction prompt
Prompt head: “Your task is to extract the entities (such as people, places, organizations, etc.)
and relations (representing behaviors or properties between entities, such as verbs,
attributes, or categories, etc.) involved in the input questions. These entities and relations can
help answer the input questions.”,
Instruction: “Please extract entities and relations in one of the following forms: entity, tuples,
or triples from the given input List. Entity means that only an entity, i.e. <entity>. Tuples means that
an entity and a relation, i.e. <entity-relation>. Triples means that complete triples, i.e. <entity-relation-entity>.
Please strictly follow the format of the example below when answering the question. ”,
Input: [The mind map $M$]
Output:
Table 8: The prompt template for extraction on global level.
Extraction prompt
Prompt head: “Your task is to extract the subgraphs involved in a set of input questions.”,
Instruction: “Please extract and organize information from a set of input questions into structured
subgraphs. Each subgraph represents a group of triples (subject, relation, object) that share common
entities and capture the logical relationships between the questions. Here are some examples:”
“ Input: [“What is the capital of France?”, “Who is the president of France?”,
“What is the population of Paris?”]
Output: [(“France”, “capital”, “Paris”), (“France, “president”, “Current President”),
(“Paris”, “population”, “Population Number”) ] ”
Input: [The mind map $M$]
Output:
Table 9: The prompt template for reasoning.
Reasoning prompt
Prompt head: “Your task is to answer the questions with the provided completed reasoning and input
knowledge.”,
Instruction: “Please note that the response must be included in square brackets [xxx].”
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Output:
Table 10: The prompt template for self-verification.
Self-verification prompt
Prompt head: “You are a logical verification assistant. Your task is to check whether the answer to a
given question is logically consistent with the provided completed reasoning and input knowledge.
If the answer is consistent, respond with “right”. If the answer is inconsistent, respond with “wrong”.”
Instruction: “Please note that the response must be included in square brackets [xxx].”
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Answer: [Answer $a^{t}_{j}$]
Output:
Table 11: The prompt template for re-thinking.
Re-thinking prompt
Prompt head:“You are a reasoning and knowledge integration assistant. Your task is to re-think
a question that was previously answered incorrectly by the self-verification model. Use the provided
completed reasoning and input knowledge to generate a new answer.”
Instruction: “Please note, if the knowledge is insufficient to answer the question, respond
with “Insufficient information, I don’t know”.The response must be included in square brackets [xxx].”
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Output: